# Adaptive Multilevel Splitting: First Application to Rare-Event Derivative Pricing

### v1

@@ Abstract
This work analyzes the computational burden of pricing binary options in rare-event settings and introduces an adaptation of the adaptive multilevel splitting (AMS) method for financial derivatives. Standard Monte Carlo is inefficient for deep out of the money binaries due to discontinuous payoffs and low exercise probabilities, requiring very large samples for accurate estimates. An AMS scheme is developed for binary options under Black–Scholes and Heston dynamics, reformulating the rare-event problem as a sequence of conditional events. Numerical experiments compare the method to Monte Carlo and to other techniques such as antithetic variables and multilevel Monte Carlo (MLMC) across four contracts: European digital calls and puts, and Asian digital calls and puts. Results show up to a 200-fold computational gain for deep out-of-the-money cases while preserving unbiasedness. No evidence is found of prior applications of AMS to financial derivatives. The approach improves pricing efficiency for rare-event contracts such as parametric insurance and catastrophe linked securities. An open-source Rcpp implementation is provided, supporting multiple discretizations and importance functions.
@@ 1 Introduction
The accurate and efficient pricing of financial derivatives is increasingly critical in modern markets, where advanced numerical methods are required for complex instruments [1]. The computational challenges of rare-event simulation extend beyond academic interest, creating bottlenecks that affect market functionality. Inaccurate pricing of low-probability events limits the ability of market makers to provide competitive quotes, reducing liquidity for these instruments [2]. This difficulty is pronounced in the insurance sector, where parametric products depend on binary triggers linked to observable parameters such as earthquake magnitude or wind speed [3, 4]. Computational limitations restrict coverage of catastrophic risks and constrain the development of innovative risk-transfer mechanisms in financial and insurance markets.
@@
These challenges are evident in binary options, which share structural similarities with parametric insurance through trigger-based payoffs. Their discontinuous structure pays a fixed amount if the underlying asset crosses a predetermined barrier at expiration and zero otherwise [5, 6]. This all-or-nothing feature makes pricing highly sensitive to the probability of rare events, particularly for deep out-of-the-money contracts where accurate tail estimation is critical.
@@
Addressing these difficulties naturally leads to simulation-based techniques. Monte Carlo methods are widely used for pricing complex derivatives due to their flexibility in high-dimensional settings [7]. The convergence rate of $O(N^{-1/2})$ creates a computational bottleneck, especially for binary options with low exercise probabilities. Reliable estimation in such cases typically requires millions of paths, rendering crude Monte Carlo impractical [8, 9].
@@
Classical variance-reduction techniques attempt to address these challenges. Antithetic variates reduce variance through negative correlation between paired samples [10, 7], but the theoretical gain is bounded by a factor of two [11]. Control variates can be more effective but require auxiliary variables that are both analytically tractable and highly correlated with the target payoff [12]. For discontinuous payoffs such as binary options, such variables are difficult to construct, limiting applicability.
@@
More advanced methods have been developed. Importance sampling modifies the probability measure to increase the frequency of rare outcomes and applies likelihood-ratio weighting to remove bias [7, 13]. Its effectiveness depends on the design of suitable distributions, which is problem-specific and difficult to generalize [14]. Another prominent approach is multilevel Monte carlo (MLMC), which reduces complexity by combining simulations on coarse and fine discretizations [15, 16]. While efficient for path-dependent derivatives, MLMC is not tailored to extreme-event pricing, focusing instead on reducing overall cost.
@@
Recent research combines these techniques to overcome individual limitations. Hybrid methods integrate MLMC with importance sampling to improve efficiency while concentrating sampling in critical regions [17, 18]. Machine learning further enhances importance sampling, with neural networks learning tilting parameters [19] and tensor-train decompositions enabling high-dimensional distribution approximation [20].
@@
This work addresses the computational challenges of binary option pricing by applying the adaptive multilevel splitting (AMS) method [21]. AMS extends classical splitting techniques for rare-event simulation [22] and builds on the foundations of sequential Monte Carlo [23]. Originally developed in reliability analysis and statistical physics [24, 25, 26, 27], AMS decomposes a rare event into a sequence of more frequent conditional events, transforming a single intractable estimation into multiple tractable subproblems. Although AMS has achieved strong results in other scientific domains, no prior applications are documented in financial derivatives pricing. Recent advances provide theoretical guarantees of unbiasedness and convergence [28, 29, 30], creating the basis for its use in finance.
@@
The contributions of this study are fourfold. First, an AMS adaptation is introduced for binary option pricing under Black–Scholes and Heston dynamics [31, 32], addressing the specific challenges of risk-neutral valuation and financial time series. Second, the sensitivity of the estimator to parameter choices, including the number of trajectories and resampling rates, is analyzed. Third, numerical experiments compare AMS to standard Monte Carlo, showing substantial gains for deep out-of-the-money options. Fourth, an open-source Rcpp implementation is released, supporting Euler, Milstein, and Andersen discretizations [33, 34, 35], two importance functions, and six binary option variants, offering a flexible toolkit for rare-event simulation in derivatives pricing.
@@
The paper is structured as follows. Section 2 reviews the background on SDE discretization, binary option pricing, and AMS methodology. Section 3 presents the limits of classical variance-reduction techniques. Section 4 illustrates the adapted AMS algorithm, establishes its theoretical properties, details the numerical implementation, and reports results against benchmark approaches. Section 5 concludes with a summary of findings and directions for future research.
@@ 2.1 Stochastic differential equation models
Numerical experiments are conducted under two standard models for asset price dynamics: the Black–Scholes model [31] and the Heston model [32]. These frameworks allow assessment of the robustness of the AMS approach across different model complexities.
For the Black–Scholes case the exact solution, obtained via logarithmic transformation, removes discretization error [36, 7]:
@@
For the Heston model the variance process requires a scheme that preserves positivity and avoids bias. The quadratic–exponential (QE) method of Andersen [35] is employed, the standard approach for accurate Heston simulation. It matches the first two conditional moments of $V_{t+\Delta t}\,|\,V_{t}$ and selects the update regime according to
@@
with $\epsilon\sim\mathcal{N}(0,1)$. The coefficients $\{K_{0},\dots,K_{4}\}$ and the parameters $a$, $b$, and $\psi$ are given explicitly in [35].
This construction preserves the positivity of variance and yields accurate joint dynamics, making it the reference scheme for Heston simulations in rare-event pricing.
@@ 2.2 Binary option pricing
Binary options are derivatives with discontinuous payoffs that depend on whether the underlying asset satisfies specific conditions. Four contracts are considered:
@@
  • asian digital put: $\text{Payoff}=\mathbf{1}_{\left\{\tfrac{1}{m}\sum_{t=1}^{m}S_{t}<K\right\}}$
The discontinuous structure makes pricing sensitive to small path variations and generates high variance in standard Monte Carlo estimates. Computational difficulties intensify for rare-event regimes, such as deep out of the money contracts, where the target probability $\mathbb{P}(A)$ is very small and required sample sizes grow inversely with its magnitude. In these settings, crude Monte Carlo becomes impractical. Binary options are therefore an effective test case for adaptive multilevel splitting: not only do they reallocate computational effort toward trajectories likely to activate the payoff condition, but their payoff naturally corresponds to the estimation of a probability, making AMS directly and rigorously applicable.
@@ 2.3 Adaptive multilevel splitting (AMS)
Adaptive multilevel splitting (AMS) [21, 24] is a variance reduction method for estimating the probability of rare events by decomposing the target set into a sequence of more probable intermediate events. Instead of brute force sampling, AMS focuses computation on trajectories that are likely to reach the rare-event region.
The idea can be illustrated with a random walk that must reach a high threshold $L_{\max}$. Rather than simulating many independent paths and counting those that succeed, AMS repeatedly removes poorly performing trajectories and replicates those that progress toward the target.
Figure 1: Illustration of the first two iterations of the AMS algorithm, where at each iteration the current threshold is $L=3$ and the worst-performing trajectory (i.e., the one with the lowest maximum) is discarded ($K=1$); a better-performing trajectory is cloned and resimulated from the time it first crossed $L$ [24]
Algorithm description: Given a Markov process $\{X_{t}\}_{t\geq 0}$ with initial distribution $\eta_{0}$, the goal is to estimate the rare-event probability $p=\mathbb{P}(X_{\tau}\in D),$ where $\tau$ is a stopping time and $D$ is the rare set. Adaptive multilevel splitting requires three key ingredients expressed here in a single narrative.
First, an importance function $\xi:\mathbb{R}^{d}\to\mathbb{R}$ measures progress toward $D$. Theorem 3.2 of [28] shows that unbiasedness holds under the relaxed condition $x\in D\implies\xi(x)\geq L_{\max}$, without the stricter equivalence $\xi(x)\geq L_{\max}\iff x\in D$. This weaker requirement is useful in financial applications, although the closer $\xi$ aligns with $D$ the lower the estimator variance.
Second, the algorithm fixes a sample size $n$ and a discard parameter $k$ with $1\leq k<n$. Choosing $1\leq k\leq n/2$ in practice maintains diversity of the trajectories.
Third, the importance of an entire trajectory $X=(X_{t})_{t\in[0,\tau_{f}]}$ is
$I(X)=\sup_{t\in[0,\tau_{f}]}\xi(X_{t}).$ (3)
The algorithm starts with $n$ i.i.d. replicas of the Markov chain $X^{0}_{j}=(X^{0}_{j,t})_{t\in\mathbb{N}}$, $j=1,\dots,n$, initialized outside $D$. At each iteration:
@@
    Compute $S^{j}=I(X^{j})$ for all replicas.
@@
    Sort $\{S^{j}\}$ and set $Z=S^{(k)}$, the $k$-th order statistic.
@@
    If $Z\geq L_{\max}$ or all $S^{j}$ are equal, terminate.
@@
    Discard the $k$ trajectories with $S^{j}\leq Z$ and replace them by clones of survivors $X^{(i)}$, $i>k$, restarted from the first crossing time of $Z$ and resimulated forward. Randomized cloning preserves unbiasedness.
  5. 5.
    Update the common weight $W\leftarrow\tfrac{n-k}{n}W$, with $W_{0}=1$.
After $Q$ iterations the probability estimator is
$\hat{p}_{\mathrm{AMS}}=W\cdot\frac{1}{n}\sum_{j=1}^{n}\mathbf{1}_{\{X^{j}\in D\}},$ (4)
which is unbiased for any admissible importance function $\xi$, and whose variance decreases as $\xi$ aligns more closely with the rare-event set.
Operatively, the adaptive multilevel splitting (AMS) algorithm proceeds as detailed in Algorithm 1:
@@
0:  Sample size $n$, discard count $k$, importance function $\xi$, final level $L_{\max}$.
@@
2:  Compute initial levels $S_{j}\leftarrow I(X_{t}^{j})$ for each trajectory $j$.
3:  Sort the levels $\{S_{j}\}_{j=1,\dots,n}$ as $S_{(1)}\leq S_{(2)}\leq\dots\leq S_{(n)}$.
4:  Set $Z\leftarrow S_{(k)}$, iteration counter $q\leftarrow 0$.
@@
6:   Determine the set of trajectories indices $J_{q}=\{j:S_{j}>Z\}$.
7:   Compute the number of trajectories to discard: $K_{q}=|\{j:S_{j}\leq Z\}|$.
8:   Discard the $K_{q}$ trajectories with $S_{j}\leq Z$.
9:   Clone exactly $K_{q}$ trajectories from the set $J_{q}$.
10:   Resimulate each cloned trajectory starting from its hitting time of the set $\{\xi>Z\}$ up to its stopping time $\tau_{j}$.
11:   Update $S_{j}\leftarrow\max_{0\leq t\leq\tau_{j}}\xi(X_{t}^{j})$ for each cloned trajectory.
12:   Sort the levels $\{S_{j}\}_{j=1,\dots,n}$ as $S_{(1)}\leq S_{(2)}\leq\dots\leq S_{(n)}$.
13:   Set $Z\leftarrow S_{(k)}$
14:   $q\leftarrow q+1$.
15:  end while
16:  Compute the final AMS estimator:
$\hat{p}_{\mathrm{AMS}}=\left(\prod_{i=0}^{q}\frac{n-k}{n}\right)\times\frac{1}{n}\sum_{j=1}^{n}\mathbf{1}_{\{X^{j}\in D\}}.$

### v2

# Adaptive Multilevel Splitting: First Application to Rare Event Derivative Pricing

@@ Abstract
This work analyzes the computational burden of pricing binary options in rare event settings and introduces an adaptation of the adaptive multilevel splitting (AMS) method for financial derivatives. Standard Monte Carlo is inefficient for deep out-of-the-money binaries due to discontinuous payoffs and low exercise probabilities, requiring very large samples for accurate estimates. An AMS scheme is developed for binary options under Black–Scholes and Heston dynamics, reformulating the rare event problem as a sequence of conditional events. Numerical experiments compare the method to Monte Carlo and to other techniques such as antithetic variables and multilevel Monte Carlo (MLMC) across three contracts: European digital calls, Asian digital calls, up-and-in barrier digital calls. Results show up to a 200-fold computational gain for deep out-of-the-money cases while preserving unbiasedness. To the best of our knowledge, this is the first application of AMS to financial derivative pricing. The approach improves pricing efficiency for rare-event contracts such as parametric insurance and catastrophe-linked securities. An open source Rcpp implementation is provided, supporting multiple discretizations and importance functions.
@@ 1 Introduction
The accurate and efficient pricing of financial derivatives is increasingly critical in modern markets, where advanced numerical methods are required for complex instruments [jrfm12010035]. The computational challenges of rare event simulation extend beyond academic interest, creating bottlenecks that affect market functionality. Inaccurate pricing of low probability events limits the ability of market makers to provide competitive quotes, reducing liquidity for these instruments [Muellerleile_2025]. This difficulty is pronounced in the insurance sector, where parametric products depend on binary triggers linked to observable parameters such as earthquake magnitude or wind speed [undici, LARSSON2023100345]. Computational limitations restrict coverage of catastrophic risks and constrain the development of innovative risk transfer mechanisms in financial and insurance markets.
@@
These challenges are evident in binary options, which share structural similarities with parametric insurance through trigger-based payoffs. Their discontinuous structure pays a fixed amount if the underlying asset crosses a predetermined barrier at expiration and zero otherwise [shreve2004stochastic2, shreve2004stochastic1]. This all-or-nothing feature makes pricing highly sensitive to the probability of rare events, particularly for deep out-of-the-money contracts where accurate tail estimation is critical.
@@
Addressing these difficulties naturally leads to simulation-based techniques. Monte Carlo methods are widely used for pricing complex derivatives due to their flexibility in high-dimensional settings [glasserman2004monte]. The convergence rate of $O(N^{-1/2})$ creates a computational bottleneck, especially for binary options with low exercise probabilities. Reliable estimation in such cases typically requires millions of paths, rendering crude Monte Carlo impractical [beck2015rareeventsimulation, bucklew2004introduction].
@@
Classical variance reduction techniques attempt to address these challenges. Antithetic variates reduce variance through negative correlation between paired samples [hammersley1956new, glasserman2004monte], but the theoretical gain is bounded by a factor of two [varredu]. Control variates can be more effective but require auxiliary variables that are both analytically tractable and highly correlated with the target payoff [rasmussen2005control]. For discontinuous payoffs such as binary options, such variables are difficult to construct, limiting applicability.
@@
More advanced methods have been developed. Importance sampling modifies the probability measure to increase the frequency of rare outcomes and applies likelihood ratio weighting to remove bias [glasserman2004monte, imps]. Its effectiveness depends on the design of suitable distributions, which is problem-specific and difficult to generalize [swiler2010importance]. Another prominent approach is multilevel Monte Carlo (MLMC), which reduces complexity by combining simulations on coarse and fine discretizations [m1, m2]. While efficient for path-dependent derivatives, MLMC is not tailored to extreme event pricing, focusing instead on reducing overall cost.
@@
Recent research combines these techniques to overcome individual limitations. Hybrid methods integrate MLMC with importance sampling to improve efficiency while concentrating sampling in critical regions [BenAlaya17022023, kebaier2018coupling]. Machine learning further enhances importance sampling, with neural networks learning tilting parameters [muller2019neural] and tensor train decompositions enabling high-dimensional distribution approximation [cui2024deep].
@@
This work addresses the computational challenges of binary option pricing by applying the adaptive multilevel splitting (AMS) method [Cérou27022007]. AMS extends classical splitting techniques for rare event simulation [garvels2000splitting] and builds on the foundations of sequential Monte Carlo [doucet2001sequential]. Originally developed in reliability analysis and statistical physics [c1, baars2021application, innes2024adaptive, refId0], AMS decomposes a rare event into a sequence of more frequent conditional events, transforming a single intractable estimation into multiple tractable subproblems. Although AMS has achieved strong results in other scientific domains, no prior applications are documented in financial derivatives pricing. Recent advances provide theoretical guarantees of unbiasedness and convergence [brehier2016unbiasedness, cerou2016fluctuation, cerou2023adaptive], creating a rigorous basis for its use in finance.
Within quantitative finance, interacting particle systems and Sequential Monte Carlo (SMC) methods have already been applied to derivative pricing, most notably in the barrier-option framework of [Shevchenko_2016]. In that setting, the option price is expressed through a Feynman–Kac representation whose potential functions enforce survival within the barrier region, and resampling is used to maintain a population of feasible trajectories. Numerical stability for rare events is typically achieved by tempering or by carefully designing these potentials, while the intermediate “levels” that guide the particle system are fixed ex ante by the time discretisation or by a prescribed sequence of thresholds. Adaptive Multilevel Splitting (AMS) can be understood as a direct evolution of this idea [c1]: instead of resampling based on fixed survival sets or generic particle weights, AMS works with a scalar importance function on the whole trajectory and introduces an adaptive sequence of levels determined by the empirical distribution of the particles. At each iteration, a fixed fraction of the least-performing particles is killed and replaced by clones of better-performing ones, so that the intermediate levels separating typical from rare trajectories are constructed on the fly.
@@
The contributions of this study are fourfold. First, an AMS adaptation is introduced for binary option pricing under Black–Scholes and Heston dynamics [due, e0f45016-f730-320a-bedc-d34f406805b2], addressing the specific challenges of risk-neutral valuation and financial time series. Second, the sensitivity of the estimator to parameter choices, including the number of trajectories and resampling rates, is analyzed. Third, numerical experiments compare AMS to standard Monte Carlo, showing substantial gains for deep out-of-the-money options. Fourth, an open source Rcpp implementation is released, supporting Euler, Milstein, and Andersen discretizations [BallyTalay1996, HighamMaoSzpruch2013, Andersen2007], two importance functions, and six binary option variants, offering a flexible toolkit for rare event simulation in derivatives pricing. A fully documented and publicly available implementation is provided in the R package amsSim [mio], which has been released on CRAN and includes all the algorithms and numerical experiments presented in this paper.
@@
The paper is structured as follows. Section 2 reviews the background on SDE discretization, binary option pricing, and AMS methodology. Section 3 presents the limits of classical variance reduction techniques. Section 4 illustrates the adapted AMS algorithm, establishes its theoretical properties, details the numerical implementation, and reports results against benchmark approaches. Section 5 concludes with a summary of findings and directions for future research.
@@ 2.1 Stochastic differential equation models
Numerical experiments are conducted under two standard models for asset price dynamics: the Black–Scholes model [due] and the Heston model [e0f45016-f730-320a-bedc-d34f406805b2]. These frameworks allow assessment of the robustness of the AMS approach across different model complexities.
For the Black–Scholes case the exact solution, obtained via logarithmic transformation, removes discretization error [BOYLE1977323, glasserman2004monte]:
@@
For the Heston model the variance process requires a scheme that preserves positivity and avoids bias. The quadratic–exponential (QE) method of Andersen [Andersen2007] is employed, the standard approach for accurate Heston simulation. It matches the first two conditional moments of $V_{t+\Delta t}\,|\,V_{t}$ and selects the update regime according to
@@
with $\epsilon\sim\mathcal{N}(0,1)$. The coefficients $\{K_{0},\dots,K_{4}\}$ and the parameters $a$, $b$, and $\psi$ are given explicitly in [Andersen2007].
This construction preserves the positivity of variance and yields accurate joint dynamics, making it the reference scheme for Heston simulations in rare event pricing.
@@ 2.2 Binary option pricing
Binary options are derivatives with discontinuous payoffs that activate when the underlying asset satisfies a prescribed condition at maturity or along the path. Three contracts are considered:
@@
  • up-and-in barrier digital call: $\text{Payoff}=\mathbf{1}_{\left\{\max_{0\leq t\leq T}S_{t}>K_{L}\right\}}$
The discontinuous structure makes pricing highly sensitive to small path variations and leads to large variance under standard Monte Carlo. Difficulties intensify in rare-event regimes, for example, deep out-of-the-money strikes, where the target probability $\mathbb{P}(A)$ is extremely small and crude Monte Carlo requires a prohibitive number of samples. These features make binary options an effective stress test for adaptive multilevel splitting: AMS reallocates computational effort toward trajectories likely to activate the payoff condition, and the payoff itself corresponds to the estimation of a probability, making the method directly and rigorously applicable.
@@ 2.3 Adaptive multilevel splitting (AMS)
Adaptive multilevel splitting (AMS) [Cérou27022007, c1] is a variance reduction method for estimating the probability of rare events. Instead of relying on brute–force Monte Carlo, which wastes almost all trajectories on paths that never approach the rare set, AMS iteratively concentrates simulation effort on trajectories that make progress towards the event of interest.
Let $\{X_{t}\}_{t\geq 0}$ be a Markov process with initial distribution $\eta_{0}$, and let
$p=\mathbb{P}\big(X_{\tau}\in D\big)$
be the probability that the process hits a rare set $D$ at a stopping time $\tau$. AMS requires three main ingredients.
Score function, rare level and trajectory score.
An importance or score function $\xi:\mathbb{R}^{d}\to\mathbb{R}$ quantifies the progress of the process towards the rare set $D$. We fix a rare level $L_{\max}\in\mathbb{R}$ and assume the relaxed condition
$x\in D\;\Rightarrow\;\xi(x)\geq L_{\max},$
which is sufficient to preserve unbiasedness [brehier2016unbiasedness]. Equality between the two sets is not required, although a score function that is more closely aligned with $D$ leads to a lower variance. For a complete trajectory $X=(X_{t})_{t\in[0,\tau_{f}]}$ we then define its score as
$I(X)\;=\;\sup_{t\in[0,\tau_{f}]}\xi(X_{t}),$ (3)
namely the maximal level reached with respect to the importance function.
Population size and killing rate.
Second, the algorithm works with a population of $N$ replicas and a killing parameter $K$ with $1\leq K<N$. At each iteration the $K$ worst replicas are removed and replaced by clones of better–performing replicas. In practice, choosing $K\leq N/2$ preserves enough diversity in the population.
The AMS mechanism can be understood by considering a process that must reach a high level $L_{\max}$.
Rather than simulating many independent trajectories from $X_{0}$ and counting only the few that cross $L_{\max}$, AMS repeatedly: (i) discards the $K$ trajectories with smallest score, and (ii) replaces them with clones of better trajectories that have already crossed an intermediate level.
The clones are restarted from the first crossing time of the current level and then resimulated onwards with fresh randomness. In this way the whole population is gradually pushed towards the rare region.
Figure 1: Illustration of the first two iterations of the AMS algorithm with population size $N=3$ and killing parameter $K=1$. The horizontal dashed lines represent successive adaptive levels $Z^{(1)}$ and $Z^{(2)}$, while the top dashed line denotes the rare level $L_{\max}$. All trajectories start from the common initial point $X_{0}$. At the first branching level, the trajectory with the lowest score is killed; one of the two better trajectories is cloned and its clone is restarted from the time at which the original path first crosses $Z^{(1)}$, after which it is resimulated using new random increments (coloured path). The same procedure is repeated at the second branching level. The figure shows how the algorithm progressively reallocates computational effort to trajectories that move closer to the rare-event level $L_{\max}$.
Algorithm description.
We now formalise the procedure underlying Figure 1. Let $X^{j}=(X_{t}^{j})_{t\in[0,\tau_{f}]}$, $j=1,\dots,N$, denote $N$ independent replicas of the Markov chain, all started outside $D$. Define their scores $S^{j}=I(X^{j})$ using (3), and let $S^{(1)}\leq\dots\leq S^{(N)}$ be the ordered scores. The algorithm maintains a weight $W$, initialised at $W_{0}=1$, and proceeds iteratively as follows:
@@
    Level selection. At iteration $q$, compute the order statistics $\{S^{(j)}\}_{j=1}^{N}$ and set
    $Z_{q}=S^{(K)},$
    the $K$-th order statistic. The level $Z_{q}$ corresponds to the horizontal dashed line labelled “$q$-th branching” in Figure 1.
@@
    Stopping criterion. If $Z_{q}\geq L_{\max}$, or if $S^{(1)}=\dots=S^{(N)}$, terminate the procedure.
@@
    Killing and cloning. Identify the index set of the $K$ worst replicas,
    $\mathcal{K}_{q}=\{j:S^{j}\leq Z_{q}\},$
    and the set of survivors $\mathcal{S}_{q}=\{1,\dots,N\}\setminus\mathcal{K}_{q}$. For each $j\in\mathcal{K}_{q}$:
    1. (a)
      Draw at random an index $i\in\mathcal{S}_{q}$.
    2. (b)
      Let
      $T_{i}(Z_{q})=\inf\{t\leq\tau_{f}:\xi(X_{t}^{i})\geq Z_{q}\}$
      be the first crossing time of the current level $Z_{q}$ by replica $i$.
    3. (c)
      Construct a new trajectory $\tilde{X}^{j}$ by setting
      $\tilde{X}_{t}^{j}=X_{t}^{i}\quad\text{for }0\leq t\leq T_{i}(Z_{q}),$
      and then simulating fresh dynamics for $t>T_{i}(Z_{q})$ starting from the state $X_{T_{i}(Z_{q})}^{i}$. In Figure 1 this corresponds to the coloured trajectory that coincides with the parent path up to the branching level and then diverges.
@@
    Weight update. Update the common weight as
    $W\;\leftarrow\;\frac{N-K}{N}\,W.$
Operatively, the adaptive multilevel splitting (AMS) algorithm proceeds as detailed in Algorithm 1.
@@
0:  Population size $N$, killing parameter $K$, importance function $\xi$, rare level $L_{\max}$.
@@
2:  Compute initial scores $S_{j}\leftarrow I(X^{j})$ for each trajectory $j$.
3:  Sort the scores $\{S_{j}\}_{j=1,\dots,N}$ as $S_{(1)}\leq S_{(2)}\leq\dots\leq S_{(N)}$.
4:  Set $Z\leftarrow S_{(K)}$, iteration counter $q\leftarrow 0$ and weight $W\leftarrow 1$.
@@
6:   Let $\mathcal{I}=\{1,\dots,N\}$ be the current index set.
7:   Define the set of candidate survivors $\mathcal{S}=\{j\in\mathcal{I}:S_{j}>Z\}$.
8:   Define the set of candidates to be killed $\mathcal{C}=\{j\in\mathcal{I}:S_{j}\leq Z\}$.
9:   Randomly choose a subset $\mathcal{K}_{q}\subseteq\mathcal{C}$ of size $K$.
10:   Set $\mathcal{S}_{q}\leftarrow\mathcal{I}\setminus\mathcal{K}_{q}$ (survivors at iteration $q$).
11:   for all $j\in\mathcal{K}_{q}$ do
12:    Randomly select a parent index $i\in\mathcal{S}_{q}$.
13:    Compute the first crossing time $T_{i}(Z)\leftarrow\inf\{t\leq\tau_{i}:\xi(X_{t}^{i})\geq Z\}$.
14:    Construct a new trajectory $\tilde{X}^{j}$ by setting $\tilde{X}_{t}^{j}\leftarrow X_{t}^{i}$ for $0\leq t\leq T_{i}(Z)$ and then resimulating forward for $t>T_{i}(Z)$ up to its stopping time $\tau_{j}$, using fresh randomness.
15:    Replace $X^{j}\leftarrow\tilde{X}^{j}$.
16:   end for
17:   Update the scores $S_{j}\leftarrow I(X^{j})$ for all $j=1,\dots,N$.
18:   Sort the scores $\{S_{j}\}$ as $S_{(1)}\leq\dots\leq S_{(N)}$ and set $Z\leftarrow S_{(K)}$.
19:   Update the weight $W\leftarrow\frac{N-K}{N}\,W$.
20:  end while
21:  Compute the final AMS estimator:
$\hat{p}_{\mathrm{AMS}}=W\times\frac{1}{N}\sum_{j=1}^{N}\mathbf{1}_{\{X^{j}\in D\}}.$